Session: Nov 2002 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Hartebeesthoek (Jonathan Quick at Sun Nov 17 07:13:01 2002) Success :
No known problems

Medicina (Fiocchi Franco at Sun Nov 17 08:06:05 2002) Minor failures : TSYS
Our tsys values are not defined because the calibration system is not able for this observation frequency.

Urumqi (Sun Zhengwen at Sun Nov 17 08:55:59 2002) Minor failures : HIGH PARI
Due to Weather,there's a little snow on the antenna surface,so the Tsys hight.

Onsala (Michael Lindqvist at Sun Nov 17 09:09:16 2002) Minor failures : RECO
1. GPS-formatter difference changed from about 6 to 56.2 us. 2. Problem with tapedrive: pass command gave pass/1,101 instead of pass/1,1

Torun (K.Borkowski at Sun Nov 17 13:15:11 2002) Success :
No known problems.

Shanghai (Huang Xinyong at Sun Nov 17 23:47:46 2002) Minor failures :
The RCP receiver is sometimes good and sometimes bad during this experiments. the data is invalid before ut04:54:00.

Effelsberg (A. Kraus at Mon Nov 18 09:23:58 2002) Success :

Jodrell_Bank (Alastair Gunn at Mon Nov 18 15:03:37 2002) Success :
No known problems.

Cambridge (Alastair Gunn at Mon Nov 18 15:04:29 2002) Success :
Due to MERLIN link restrictions, signal only appears in BBCs 3 and 4. No known problems. No valid Tsys measurements.

Arecibo (Tapasi Ghosh at Wed Nov 20 14:39:06 2002) Minor failures : LATE RECO
1. One of the write drivers was removed from the recorder to investigate the "Track 18 problem". As it wasn't possible to place it back before the run, we went ahead and ran the experiment in the usual way without the write-driver # 2. Hence, headstack 2 didn't write any data on tape (please accept our apologies. Meanwhile, Track #18 trouble has been located to a loose capacitor within that module and it is beleived to be fixed now). 2. As the one of the local control computers needed rebooting, the first 4 mins of the run couldn't be recorded.

Westerbork (Tony Foley at Thu Dec 12 12:47:43 2002) Minor failures : PHAS
VC3/VC4, VC7/VC8 swap